# EMOD: A Unified EEG Emotion Representation Framework Leveraging V-A Guided Contrastive Learning

### v1

@@ Conclusion
In this work, we proposed EMOD, a contrastive pretraining framework for EEG-based emotion recognition. EMOD unifies heterogeneous emotion labels via Valence–Arousal guidance and employs a Triple-Domain Encoder with a Spatial-Temporal Transformer to model diverse EEG features. Extensive experiments on three benchmarks demonstrate that EMOD outperforms existing methods while using fewer parameters and less pretraining data. Ablation studies further validate the importance of our emotion-aware contrastive loss and architectural design. Our findings highlight the potential of emotionally guided pretraining for learning generalizable and affect-sensitive EEG representations.

### v2

@@ Conclusion
In this work, we proposed EMOD, a contrastive pretraining framework for EEG-based emotion recognition. EMOD unifies heterogeneous emotion labels via Valence-Arousal guidance and employs a Triple-Domain Encoder with a Spatial-Temporal Transformer to model diverse EEG features. Extensive experiments on three benchmarks demonstrate that EMOD outperforms existing methods while using fewer parameters and less pretraining data. Ablation studies further validate the importance of our emotion-aware contrastive loss and architectural design. Our findings highlight the potential of emotionally guided pretraining for learning generalizable and affect-sensitive EEG representations.
Acknowledgments
This work was supported by STI 2030 Major Projects (2021ZD0200400), the National Natural Science Foundation of China (No.62476240), the Fundamental Research Funds for the Central Universities (226-2025-00122) and the Key Program of the Natural Science Foundation of Zhejiang Province, China (No. LZ24F020004). The corresponding author is Dr. Sha Zhao.